# Tiny Time Mixers (TTMs): Fast Pre-trained Models for Enhanced Zero/Few-Shot Forecasting of Multivariate Time Series

### v6

@@ 4.3 TTM Model Details
Although, a TTM model needs to be pre-trained for a specific forecast length (FL), we provide two forecast length adaption (FLA) techniques (explained in Section 5) that enable a pre-trained TTM to work across different FLs. Users can either build a direct pre-trained model (from one of the above variants) targeting a specific FL, or use the FLA techniques to adapt an existing TTM model to their application setting. Primary results are reported using the direct approach, and a detailed ablation study is provided to compare the effectiveness of various FLA techniques. In the direct approach, model parameter size varies across FLs and we report the average parameter size in the result tables. TTM fine-tuning and inferencing are highly efficient and fast, requiring only 1 GPU or even CPU execution. All model hyperparameters are chosen based on validation performance, and final test results are reported. Refer to Appendix D for detailed model specifications and hyper-parameters.

### v7

@@ 4.3 TTM Model Details
Although, a TTM model needs to be pre-trained for a specific forecast length (FL), we provide two forecast length adaption (FLA) techniques (explained in Section 4.7) that enable a pre-trained TTM to work across different FLs. Users can either build a direct pre-trained model (from one of the above variants) targeting a specific FL, or use the FLA techniques to adapt an existing TTM model to their application setting. Primary results are reported using the direct approach, and a detailed ablation study is provided to compare the effectiveness of various FLA techniques. In the direct approach, model parameter size varies across FLs and we report the average parameter size in the result tables. TTM fine-tuning and inferencing are highly efficient and fast, requiring only 1 GPU or even CPU execution. All model hyperparameters are chosen based on validation performance, and final test results are reported. Refer to Appendix D for detailed model specifications and hyper-parameters.